OpenStudy (anonymous):

Whats the solution of these two equations? 2y=8-x y-2x= -1

OpenStudy (amistre64):

there are a few methods that could be implemented to find a solution if it exists

OpenStudy (anonymous):

system of equations allegbricially

OpenStudy (anonymous):

@Helllo;3 do you know how to solve for y in the first equation?

OpenStudy (amistre64):

yes, thats tends to cover a few methods tho

OpenStudy (anonymous):

noo , and alsoo i need bothh

OpenStudy (anonymous):

Solve for y. Just get y on the left side and everything else on the right. Doesn't matter if there's an x. I'll guide you through this.

OpenStudy (anonymous):

like solution ( ) , ( ) the corindatinates

OpenStudy (anonymous):

okaay

OpenStudy (anonymous):

What do you do to solve for y?

OpenStudy (anonymous):

you subtract and do the inverse

OpenStudy (anonymous):

2y=8-x You have 2*y. What do you need to do to get that 2 on the other side?

OpenStudy (anonymous):

can you draww it , V.V

OpenStudy (anonymous):

you substract 2 from both sides

OpenStudy (amistre64):

eq2 might be simpler to solve for y :)

OpenStudy (anonymous):

2y/2 = (8-x)/2

OpenStudy (anonymous):

l need the solution for both x nd y

OpenStudy (amistre64):

helllo, if your just expecting an answer then you really not going to like what we have to offer. if you work thru it with briney youll do better

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Yes. And it is a step-by-step process. You have to solve for one variable first. In this case, you get the y on one side and the x along with other terms on the other side. y = some stuff plug that into the second equation for y. y-2x= -1 (some stuff) -2x =-1 You solve for x. You get: x= a number. You plug that back into the first equation to get y. 2y=8-x 2(a number) = 8-x Then you solve for x. Then you're done.

OpenStudy (anonymous):

what you meaan im done

OpenStudy (anonymous):

Those are the steps to solve in a systems of equations. It's not just a one-step process.

OpenStudy (amistre64):

like alice in wonderland; you start at the begining, and when you get to the end ... your done lol

OpenStudy (anonymous):

im sorry , but i really dont understand this ,

OpenStudy (anonymous):

lol ,

OpenStudy (anonymous):

I can guide you through to getting the answer to this one. But after that, throw another system of equations out there and see if you can solve it given the process I just told you.

OpenStudy (anonymous):

okayy :)

OpenStudy (amistre64):

in order to be good at systems of equations you have to have a firm grasp on how to manipulate a single equation to solve for any particular variable that arises in it

OpenStudy (amistre64):

step one, find an equation that would solve for y from one of the 2 that are offerec

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

what is offerec

OpenStudy (amistre64):

yeah, you reallyhave to read thru the typing mistakes

OpenStudy (amistre64):

offered

OpenStudy (anonymous):

okay soo list the steps for mee ,

OpenStudy (amistre64):

briney listed them

OpenStudy (anonymous):

soo i know what to do

OpenStudy (anonymous):

im lostt ,,,

OpenStudy (amistre64):

me too, this little screen aint good for the site

OpenStudy (amistre64):

i see about 2 lines of text :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

draw it for me pleasee

OpenStudy (anonymous):

2y=8-x y-2x= -1 1. Solve for y. 2y = 8-x To get rid of the two, you have to DIVIDE by 2 on each side. (2y)/2 = (8-x)/2 y = 8/2 -x/2 y=4 - x/2 2. Plug what you got for y into the "y" in your second equation. y-2x= -1 (4 - x/2) - 2x = -1 4 -x/2 -2x = -1 -x/2 -4x/2 = -5 -5x/2 = -5 (-2/5)*(-5x/2) = -5(-2/5) x = 2 3. Plug x=2 into eqn 1. 2y=8-x 2y = 8 -(2) 2y = 6 (2y)/2 = 6/2 y=3 Okay I am sure this is correct. I missed a "-1" in one of my steps. But there you go. 3 steps to finding your solutions. x=2 and y=3

OpenStudy (amistre64):

using a usb keyboard makes the typing easier ...

OpenStudy (anonymous):

ohh , nd can yu on britney draw it ,pleasee

OpenStudy (anonymous):

I don't have a drawing pad. And if you see my writing in real life, it's bad. On a drawing board with no writing pad = illegible.

OpenStudy (anonymous):

I think it would help you a lot more if you watch a youtube vid. I'll find one. If you study the 3-step method and do a ton of problems, it will become really easy.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Along with watching the video, I would suggest writing everything I typed out on a piece of paper. Writing things out helps you learn the material.

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

He really gets into solving these at 1:46 http://www.youtube.com/watch?v=cwHR_B9zK7k

OpenStudy (anonymous):

okay let me see if it will hlp

OpenStudy (anonymous):

help

OpenStudy (anonymous):

ooo i seee ,

OpenStudy (anonymous):

okay now im going to try myself

OpenStudy (anonymous):

Yay! Yeah, PatrickJMT has some really good videos. Paul's online notes are good too: http://tutorial.math.lamar.edu/Classes/Alg/SystemsTwoVrble.aspx